OpenStudy (anonymous):

may someone help with this.. 1/9x = 6 = 1/3x

OpenStudy (anonymous):

must be a typo with the two equal signs

OpenStudy (anonymous):

yes sorry the first equal sign should be a plus sign

OpenStudy (anonymous):

\[\frac{1}{9}x+6=\frac{1}{3}x\]??

OpenStudy (anonymous):

yes thats the equation

OpenStudy (anonymous):

multiply both sides by \(9\)

OpenStudy (anonymous):

oh ok thats simple thank you!

OpenStudy (anonymous):

\[x+54=3x\] and continue form there

OpenStudy (anonymous):

ok thanks!